Solve using substitution. Solve first equation to get the y and substitute the answer into the deco d equation to find the x. X-6y=11 and 2x-5y=1
Answer by stanbon(75887)

Solve using substitution. Solve first equation to get the y and substitute the answer into the deco d equation to find the x.
X-6y=11 and 2x-5y=1
----
y = (x-11)/6
------
2x -5(x-11)/6) = 1
-----
12x -5x + 55 = 6
--------
7x = -49
-------
x = -7
=============
Cheers,
Stan H.
=================